# LaTo: Landmark-tokenized Diffusion Transformer for Fine-grained Human Face Editing

### v1

@@ Abstract
Recent multimodal models for instruction-based face editing enable semantic manipulation but still struggle with precise attribute control and identity preservation. Structural facial representations such as landmarks are effective for intermediate supervision, yet most existing methods treat them as rigid geometric constraints, which can degrade identity when conditional landmarks deviate significantly from the source (e.g., large expression or pose changes, inaccurate landmark estimates). To address these limitations, we propose LaTo, a landmark-tokenized diffusion transformer for fine-grained, identity-preserving face editing. Our key innovations include: (1) a landmark tokenizer that directly quantizes raw landmark coordinates into discrete facial tokens, obviating the need for dense pixel-wise correspondence; (2) a location-mapping positional encoding that integrates facial and image tokens for unified processing, enabling flexible yet decoupled geometry-appearance interactions with high efficiency and strong identity preservation; and (3) a landmark predictor that leverages vision–language models to infer target landmarks from instructions and source images, whose structured chain-of-thought improves estimation accuracy and interactive control. To mitigate data scarcity, we curate HFL-150K, to our knowledge the largest benchmark for this task, containing over 150K real face pairs with fine-grained instructions. Extensive experiments show that LaTo outperforms state-of-the-art methods by 7.8% in identity preservation and 4.6% in semantic consistency. Code and dataset will be made publicly available upon acceptance.
@@ 1 Introduction
Figure 1: Landmark tokenization in LaTo preserves identity and produces natural results, whereas pixelwise alignment baselines rigidly follow the rendered landmark image and often lose identity under substantial differences in pose, expression, or face shape.
A common strategy for improving edit fidelity is to employ facial landmarks as an intermediate structural prior (Yang & Guo, 2020; Wei et al., 2025; Liang et al., 2024). Unlike text prompts, landmarks (Li et al., 2022; Sun et al., 2024) impose explicit geometric constraints via precise 2D coordinates of key facial features (eyes, nose, mouth), thereby localizing edits to the appropriate regions. However, most existing approaches are built on GANs (Goodfellow et al., 2020) or UNet-based (Ronneberger et al., 2015) diffusion models and transfer poorly to modern Diffusion Transformer (DiT) (Peebles & Xie, 2023) due to fundamental architectural differences. Recent DiT-based editors like OminiControl (Tan et al., 2024) and OmniGen (Xiao et al., 2025) adopt a general-purpose control strategy for face editing: they rasterize landmarks into 2D images, encode them via Variational Autoencoder (VAE) (Kingma & Welling, 2022) to obtain dense visual tokens, and use these as in-context guidance. Despite improving identity preservation, this strategy introduces two core limitations: (1) conditioning on rendered landmark images encourages pixel-wise copying of fixed facial shapes rather than geometric reasoning, leading to identity drift and artifacts when the conditional landmarks substantially deviate from the source in shape or position, as shown in Figure 1; and (2) because self-attention scales quadratically with sequence length (Huang et al., 2024; Avrahami et al., 2025), appending long dense visual tokens to diffusion tokens incurs prohibitive memory and compute costs, limiting practical applicability in complex scenarios.
@@ 3.1 HFL-150K Dataset Construction
Large-scale editing datasets have been proven critical for developing advanced editing models. In face editing, existing datasets such as SeqDeepFake (Shao et al., 2022) and SEED  (Zhu et al., 2025) suffer from two fundamental limitations: (1) they rely on coarse-grained facial attribute instructions and outdated synthesis models (Karras et al., 2019; Tsaban & Passos, 2023), resulting in unrealistic editing artifacts and limited result diversity; (2) their scale is constrained by poor-quality samples that require extensive filtering to remove invalid examples. To address the limitations of existing benchmarks, we introduce HFL-150K, a large-scale human face editing dataset comprising 150k image-edit instruction triplets (source, instruction, edited). As summarized in Table 1, HFL-150K is constructed through a hybrid approach combining real-world image curation and synthetic generation using advanced editing models.
@@ 3.2.1 Landmark Tokenizer
Figure 4: Overview of LaTo. The landmark predictor infers target landmarks from source image and instruction via structured chain of thought. A landmark tokenizer and visual VAE encode predicted landmarks and source image into tokens. A token fuser merges tokens with stochastic noise and feeds them to DiT blocks, whose denoising yields a target image aligned with landmarks and instruction.
@@ 4.3 Ablation study
rendered image
visual VAE / 0.821 0.712 0.744 0.584 83.6 1.76
@@
coordinates
landmark
tokenizer
(1) w/o PE 0.654 0.611 0.630 0.512 50.7 58.9
@@
[…]
Landmark-aware CFG scaling analysis. Figure 6 and Table 4.3 illustrate the impact of our CFG scale. Raising the CFG scale improves landmark alignment, yet may generate more artefacts and compromise video quality. We adopt a CFG scale of 4 as the optimal baseline configuration.
Figure 6: Visualization of the landmark-aware CFG scale $w$.
@@ C.1 Scaling Analysis of HFL-150K
C.2 Comparison of Landmark Predictors
@@
C.3 Design Choices for Unconditional Landmark Tokens
@@
C.4 Rectified Identity Preservation Score
@@
C.5 Extended Qualitative Comparisons

### v2

@@ Abstract
Recent multimodal models for instruction-based face editing enable semantic manipulation but still struggle with precise attribute control and identity preservation. Structural facial representations such as landmarks are effective for intermediate supervision, yet most existing methods treat them as rigid geometric constraints, which can degrade identity when conditional landmarks deviate significantly from the source (e.g., large expression or pose changes, inaccurate landmark estimates). To address these limitations, we propose LaTo, a landmark-tokenized diffusion transformer for fine-grained, identity-preserving face editing. Our key innovations include: (1) a landmark tokenizer that directly quantizes raw landmark coordinates into discrete facial tokens, obviating the need for dense pixel-wise correspondence; (2) a location-mapped positional encoding and a landmark-aware classifier-free guidance that jointly facilitate flexible yet decoupled interactions among instruction, geometry, and appearance, enabling strong identity preservation; and (3) a landmark predictor that leverages vision–language models to infer target landmarks from instructions and source images, whose structured chain-of-thought improves estimation accuracy and interactive control. To mitigate data scarcity, we curate HFL-150K, to our knowledge the largest benchmark for this task, containing over 150K real face pairs with fine-grained instructions. Extensive experiments show that LaTo outperforms state-of-the-art methods by 7.8% in identity preservation and 4.6% in semantic consistency. Code and dataset will be made publicly available upon acceptance.
@@ 1 Introduction
Figure 1: Landmark tokenization in LaTo preserves identity and produces natural results, whereas pixelwise alignment baselines rigidly follow the rendered landmark image and often lose identity under cross-identity landmark conditions (first four columns) or when self-identity landmarks differ substantially from the source.
A common strategy for improving edit fidelity is to employ facial landmarks as an intermediate structural prior (Yang and Guo, 2020; Wei et al., 2025; Liang et al., 2024). Unlike text prompts, landmarks (Li et al., 2022; Sun et al., 2024) impose explicit geometric constraints via precise 2D coordinates of key facial features (eyes, nose, mouth), thereby localizing edits to the appropriate regions. However, most existing approaches are built on GANs (Goodfellow et al., 2020) or UNet-based (Ronneberger et al., 2015) diffusion models and transfer poorly to modern Diffusion Transformer (DiT) (Peebles and Xie, 2023) due to fundamental architectural differences. Recent DiT-based editors like OminiControl (Tan et al., 2024) and OmniGen (Xiao et al., 2025) adopt a general-purpose control strategy for face editing: they rasterize landmarks into 2D images, encode them via Variational Autoencoder (VAE) (Kingma and Welling, 2022) to obtain dense visual tokens, and use these as in-context guidance. Despite improving identity preservation, this strategy introduces two core limitations: (1) conditioning on rendered landmark images encourages pixel-wise copying of fixed facial shapes rather than geometric reasoning, leading to identity drift and artifacts when the conditional landmarks substantially deviate from the source in shape or position, as shown in Figure 1; and (2) because self-attention scales quadratically with sequence length (Huang et al., 2024; Avrahami et al., 2025), appending long dense visual tokens to diffusion tokens incurs prohibitive memory and compute costs, limiting practical applicability in complex scenarios.
@@ 3.1 HFL-150K Dataset Construction
Large-scale editing datasets have been proven critical for developing advanced editing models. In face editing, existing datasets such as SeqDeepFake (Shao et al., 2022) and SEED  (Zhu et al., 2025) suffer from two fundamental limitations: (1) they rely on coarse-grained facial attribute instructions and outdated synthesis models (Karras et al., 2019; Tsaban and Passos, 2023), resulting in unrealistic editing artifacts and limited result diversity; (2) their scale is constrained by poor-quality samples that require extensive filtering to remove invalid examples. To address the limitations of existing benchmarks, we introduce HFL-150K, a large-scale human face editing dataset comprising 150k image-edit instruction triplets (source, instruction, edited). As summarized in Table 1, HFL-150K is constructed through a hybrid approach combining real-world image curation and synthetic generation using advanced editing models.
@@ 3.2.1 Landmark Tokenizer
Figure 4: Overview of LaTo. The landmark predictor infers target landmarks from source image and instruction via structured chain of thought. A landmark tokenizer and visual VAE encode predicted landmarks and source image into tokens. The location-mapping positional encoding anchors each landmark token to its physical location, ensuring unified yet flexible alignment with instruction and visual tokens. The learned unconditional landmark token further guides the denoising process, keeping the edited image aligned with both the specified landmarks and instructions.
@@ 4.3 Ablation study
rendered image visual VAE / 0.821 0.712 0.744 0.584 83.6 1.76
@@
coordinates landmark tokenizer (1) w/o PE 0.654 0.611 0.630 0.512 50.7 58.9
@@
Figure 6: Qualitative results on challenging inputs, including cross-identity landmarks, incomplete landmarks, and stylistic inputs. For cross-identity cases, the corresponding driving images are shown in the bottom-right corner.
@@
[…]
Landmark-aware CFG scaling analysis. Figure 8 and Table 4.3 illustrate the impact of our CFG scale. Raising the CFG scale improves landmark alignment, yet may generate more artifacts and compromise video quality. We adopt a CFG scale of 4 as the optimal baseline configuration.
[…]
Figure 7: Visualization of the performance of landmark predictor under ambiguous instructions. Predicted landmarks are overlaid on the source images.
Figure 8: Visualization of the landmark-aware CFG scale $w$.
@@ 5 Conclusion
6 Acknowledgements
This work is supported by National Natural Science Foundation of China (62571322, 62431015, 62271308), STCSM (24ZR1432000, 24511106902, 24511106900, 22DZ2229005), 111 plan (BP0719010), and State Key Laboratory of UHD Video and Audio Production and Presentation.
@@ C.1 Scaling Analysis of HFL-150K
C.2 User study
[…]
Table 8: User study results.
Methods SC VQ IP Overall
Step1X-Edit† 0.728 0.810 0.746 0.761
Flux.1-Kontext† 0.712 0.824 0.774 0.770
LaTo 0.766 0.837 0.829 0.811
C.3 Extension to Video-based Facial Animation
We extend both the Landmark Tokenizer and the location‑mapping positional encoding to facial animation, enabling controllable video‑avatar generation. The key differences from LaTo are as follows. For data, we construct a 16k high‑quality single‑portrait dataset from OpenHumanVid using similar filtering criteria, with each video annotated by landmark trajectories and descriptive captions. For the tokenizer, we generalize it to 3D with a temporal length of 81 frames. For fusion, we adopt Wan2.2‑TI2V‑5B (Wan et al., 2025) as the base model and introduce an additional cross‑attention layer after the text‑to‑video attention. During training, we freeze all original weights and only train this new cross‑attention, which makes the adaptation efficient and avoids disrupting the pretrained text‑video alignment. The visual results in Figure 11 show that LaTo maintains robust identity preservation when editing with cross‑identity reference face videos, demonstrating strong scalability. However, the current system is reference‑video driven, and further work is needed to support interactive control similar to LaTo. These include developing a reliable landmark predictor for generating landmark sequences and identifying a more optimal fusion strategy, as the current additional cross‑attention may be suboptimal.
Figure 11: Examples of facial animation generated by LaTo. Given a cross‑identity driving video, LaTo can also achieve identity–geometry decoupling in the video domain, preserving the target identity while following the facial motions of the driving video.
C.4 Comparison of Landmark Predictors
@@
C.5 Design Choices for Unconditional Landmark Tokens
@@
C.6 Rectified Identity Preservation Score
@@
C.7 Extended Qualitative Comparisons
@@ Appendix D Additional Methodological Details
We also evaluated codebook sizes of 1,024, 8,192, and 16,384, as well as a no‑quantization ablation. Table 9 reports landmark reconstruction performance on HFL‑150K. For PSNR evaluation, we render the landmarks into images. Removing quantization forces the tokenizer to operate on continuous and inconsistent facial features, which degrades performance. This confirms that discrete and general landmark tokens provide more stable geometric representations and improve generalization. Although larger codebooks offer higher nominal capacity, the 16,384‑entry codebook resulted in unstable optimization and significant codebook underutilization, which reduced its effective capacity. The 8,192‑entry codebook achieved the best overall performance. We anticipate further gains through improved codebook utilization.
Table 9: Ablation of codebook size.
Codebook Size Utilization (%) L1 Distance PSNR
no-quantization - 7.04 19.11
1024 81.42 5.32 20.47
8192 70.39 3.67 21.33
16384 44.75 4.51 20.61
@@ 4. Edited Landmarks JSON:
D.3 Face Editing Questionnaire
Thank you for completing the following rating of image editing results. Please spend about 15 seconds on each image, consider the source image and the editing instruction, and rate the result based on the following dimensions:
1. Semantic Consistency
Does it match the text prompt (e.g., “Smiling slightly and turn her head left 30 degrees”)?
1: No match at all.
2: Very weak or incorrect interpretation (e.g., wrong expression or pose direction).
3: Partial match — correct type but inaccurate intensity or rotation degree.
4: Mostly accurate — small deviation in expression strength or head angle.
5: Perfectly matches the prompt in both expression and head pose.
2. Visual Quality
How photorealistic and artifact-free is the generated face?
1: Severely distorted or unrealistic — obvious artifacts, broken facial structure, or non-human appearance.
2: Low quality — blurry, noisy, or with noticeable distortions (e.g., warped eyes, asymmetric features).
3: Acceptable but flawed — overall plausible, but has mild blurriness, texture glitches, or minor asymmetry.
4: High quality — sharp and realistic, with only subtle imperfections hard to notice at a glance.
5: Excellent quality — indistinguishable from a real photo; no visible artifacts or distortions.
3. Identity Preservation
Is it the same person as the source?
1: Completely different person.
2: Unlikely the same — key facial features (e.g., eye spacing, nose shape, jawline) don’t match.
3: Uncertain — could be the same or different.
4: Likely the same — minor structural differences, but core identity preserved.
5: Definitely the same person — identity is unmistakable.
Appendix E Limitations
Though LaTo is effective for fine-grained editing and identity preservation, the complex architecture of the advanced base model limits its ability to support efficient on-the-fly processing, so we rely on future advances in accelerating such base models by the community. Moreover, we observe that for input images with extreme head poses caused by heavy occlusion, landmark estimation becomes less reliable, leading to slight degradation in identity preservation. We plan to incorporate 3D-aware priors and richer occlusion annotations to mitigate these limitations.